OpenStudy (anonymous):

Pleasee help! 1/4(x-2)-1/5(x-3)=1/2 HOW do i figure this out?!

OpenStudy (phi):

is this \[ \frac{1}{4(x-2)}-\frac{1}{5(x-3)}= \frac{1}{2} \] ?

OpenStudy (anonymous):

nooo whenever there's a / its a fraction.

OpenStudy (anonymous):

 sorry its messy.